This is NA61/SHINE electronic logbook (read-only BACKUP!)

Run summary Shift logbook Run settings Beam logbook BPD logbook DAQ logbook DCS logbook Experts logbook Gas logbook TOF logbook Trigger logbook PSD logbook VD logbook Run coordination Detector Upgrade DRS logbook
  Logbook for all experts' activities  NA61/SHINE logo
Message ID: 135     Entry time: Tue, 10.07.2012 10:41:14
Author: A.Laszlo 
Subject: Magnet settings for 2012 158GeV/c p+Pb data taking 
List of responsables in 2012:
- Magnet cryogenics: Laura Stuart - 160720
- Magnet power supply: Yves Gaillard - 160950, Xavier Genillon - 162019, on call service - 163668
- Magnet: Alexei Dudarev - 164016

Magnet history in 2012:
- Problem with VTX2 cryogenics, solved.
- Problem with VTX1 and VTX2 current lead cooling water: flooded the area, especially VTX2 downstream side due to unterminated water pipe. Solved: cut water connections, closed valves. These lines are not necessary (were never used before).
- Problem with ramping up VTX1. First around 2000A some 40mV amplitude and 10ms duration spikes appeared on the voltage difference of the up and down coils of VTX1. Solved: added higher voltage alarm limit to magnet software with 100ms time tolerance window. Then, around 4000A the cryogenics were not regulated in proportion to current. Solved: software bug in cryogenics software.
- Problem with reading exact currents on GUI display. Temporary workaround: click "Terminal Window", then type "adc16 2". Software people came to solve, now it seems OK.
- The short small amplitude spikes of VTX1 are not observed anymore after proper fixing of S4.
- 2012.07.19 around 07:00: VTX1 loses vacuum and quenches.

Magnet settings in 2012:
- For 158GeV/c p beam: VTX1 (set: 5011.6A, read: -5011.89A), VTX2 (set: 3774.4A, read: -3779.26A). Reference value from 23.08.2000: VTX1 (read: -5011.6A), VTX2 (read: -3779.5A).
ELOG V3.1.3-7933898